|Version 1 (modified by 16 months ago) ( diff ),|
APO 3.5-m Users Committee Phonecon, 2/06/17
Attending: Bill Ketzeback, Nancy Chanover, Jon Holtzman, John Bally, Rachel Kuzio de Naray, Mary Beth Kaiser, Joanne Hughes Clark, John Tobin, Anne Verbiscer, Scott Anderson, Ben Williams, Sarah Tuttle, Jason Lozo, and Chip Kolbulnicky
User feedback and comments from institutional representatives:
- Washington - Scott had nothing new to report. Regarding a subsequent agenda item, he has had no comments or interest yet from his users regarding WIRO time trades.
- NMSU - Jon mentioned some broken filter trace data links on the APO web page; he is forwarding the information to Bill.
- Colorado - John B. had nothing new to report.
- JHU - Mary Beth reported that a JHU user was unable to access the APO wiki and wanted to ask more about it. Jon Holtzman stated that it is hosted by NMSU Astronomy and it is open to view by everyone, however, to post something to the wiki a user needs to create an account with a password. Jon will investigate the JHU problem further.
- Virginia - Anne had nothing to report, but queried about an ARCTIC reduction cookbook. Nancy explained that there was a link to one of Joe Huehnerhoff's webpages on the wiki but that link broke when Joe's UW account went away. We have since located the document and she will post it to the wiki. There are a number of python scripts as well that Rachel explained are still linked to for quad processing of ARCTIC data that point to Joe's github page. We likely will need to move those.
- Georgia State - Rachel had nothing new to report.
- Oklahoma - John T. had nothing new to report.
- Wyoming - Chip had nothing new to report.
- Seattle - Joanne had nothing new to report. Re: the ARCTIC reduction tools discussion, she mentioned that Karen Kinemuchi had sent her some QUADPROC (IRAF) data to review and she will get to that soon.
- NAPG - No report.
- Adler - No report.
Telescope and Instruments Report (Ketzeback):
Bill addressed the two telescope problems mentioned in the highlights section. He explained that these are both known problems for which we have work-arounds, but which would be quite involved to fix.
Bill mentioned that we started warming up TripleSpec over the weekend and that the instrument will be unavailable for several weeks. This is a step in the troubleshooting procedure we have worked out with the UVa instrument builders, although it is not clear whether this will correct the instrument problem. If it does not, the instrument will be opend up to investigate, likely over the summer, and in that case would not be available again until after summer shutdown.
The ARCTIC diffuser had a problem last night. It is unclear why it stopped rotating but is working now. It was tested without issue for 30 minutes this morning.
Mark’s detailed report follows below.
3.5-m Telescope and Instruments Highlights, 1/5/17 through 2/1/17
Several winter storms came through back to back leaving about a foot of snow. Despite very high winds we only had a couple power outages and no problems as site stayed on power (minus a transfer switch fault for the shop). We did have loss of air pressure one day but no ill effects from it. Due in part to storms, observing time has been minimal except for last several days.
1) Telescope - Telescope was operational with one issue on M2 control. This turned out to not be a new issue and occurs when large moves are requested first thing after sitting for a while. Hardware fix is very involved but we have made a few software and procedural changes to mitigate the effects of this. We also had one telescope freeze up at 60 deg during the end of night. Again problem is understood but fix is involved.
DIS blue and red cameras have been operational. The red camera has showed some indication of return of contamination but still at a minimal amount. A Conceptual PDR for several possible upgrades was held and covered later.
TripleSpec has produced only corrupted images during this last period. We are planning for a temperature cycle in the next 30 days and if that doesn’t work then a UVa servicing mission in the summer. For all practical matters TSpec is not available.
Agile has been operational this period.
Echelle has been operational with no problems.
NIC-FPS is operational.
ARCTIC is operational. The reduction in thermal strapping and consequentially slightly warmer operating temperature seems to have helped with loss quadrant issue. The diffuser now has a precision bearing and is operating as required. We are still looking at a different motor option but for now diffuser is working properly.
3) ARCSAT - Has been performing nominally. Occasional slow downs are corrected by obs spec restarting systems.
Q1 3.5m schedule updates:
Ben reported that Russet recently updated the online version of the schedule. He asked people to make certain they had the most recent version. Otherwise, nothing to report for this quarter.
Q2 3.5m scheduling:
Ben has sent out time allocations to each of the institutional schedulers. He also reminded users that we are now using a slightly revised form for proposing for time. There is a new section for ToO requests.
Report on DIS upgrades PDR
Nancy gave a brief overview of who attended and what topics were discussed at the PDR. A clear set of action items was identified during the meeting. Sarah Tuttle is generating a summary report document that will be shared with users soon. Nancy felt the review was very useful in identifying all the issues and potential paths forward. There was a brief discussion of grating changes. We would like to go back to normal grating changes, however, we have developed a plan for further engineering monitoring of environmental conditions inside the instrument grating bay that we wish to further understand first. We wish to be very cautious about grating changes during periods of high humidity (> 60%), for instance. We also plan to implement a more frequent monitoring program to better understand the scattered light / contamination issue. If users do have a pressing need for alternative gratings for their science they should let us know and we will take those under consideration when we are doing the scheduling.
open from previous meetings:
- User committee members to poll their users to see if other new slit configurations are needed at the telescope, and get back to Nancy by 19 January.
Status: Closed. Nancy received no additional input or requests from users.
- Committee members to query their users on their interest in using the WIRO wide field imaging camera (see: http://physics.uwyo.edu/~WIRO/DoublePrime/doubleprime.html). Seattle U, CU, and UVa expressed interest. (NAPG did too, by email). Nancy will discuss potential trades with these people offline.
new from this meeting:
- Nancy will work with Jason, Rachel, Jon H. and others to get the ARCTIC information on the APO wiki cleaned up and broadly accessible to ARC users.
The next Users Committee phonecon will be on Monday 13 March at 1030 MST — the agenda and other materials will be sent to committee members during the preceding week. NOTE: we may switch our telecons to the second Monday of every month to avoid a schedule conflict for one of the members. We will try it for March and see how it goes.